eliminate the arbitrary constant using any method

12.y\:=\:3Ax^3\:+\:2Bx^2\:+\:4Cx\:−\:D


1
Expert's answer
2021-09-07T03:04:37-0400

"y=3Ax^3+2Bx^2+4Cx-D
\\\frac{dy}{dx}=9Ax^2+4Bx+4C ...(1)
\\\frac{d^2y}{dx^2}=18Ax+4B ...(2)
\\\frac{d^3y}{dx^3}=18A ...(3)
\\\frac{d^4y}{dx^4}=0"